OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (mertsj):

Does 7^2+24^2=25^2?

OpenStudy (hoblos):

25^2 = 625 7^2 + 24^2 = 625 so the side lengths form a Pythagorean triple since 25^2 = 7^2+24^2,

OpenStudy (anonymous):

for a Pythagorean triplet, sum of squares of the two smaller numbers has to be equal to the the square of the third number

OpenStudy (anonymous):

thanks all of you guys. You're geniuses. It's crazy how you can just answer stuff like this off the top of your head. I'll forget it by 11th grade haha.

OpenStudy (anonymous):

As hoblos has shown, they form a Pythagorean Triplet

OpenStudy (anonymous):

Mhm, thanks c: